# Characterization of Gaussian Tensor Ensembles

### v1

@@
Abstract
The starting point of this work is a theorem due to Maxwell characterizing the distribution of a Gaussian vector with at least two coordinates. We define the Gaussian Orthogonal, Unitary and Symplectic Tensor Ensembles for notions of real symmetric, hermitian and self-dual hermitian tensors which recover the classical vector and matrix Gaussian Ensembles when the order is one and two. We give a basis of invariant polynomials for orthogonal, unitary and symplectic transformations and we prove a Maxwell-type theorem for these Gaussian tensor distributions unifying and extending the ones known for vectors and matrices.
@@ 1.1 Characterization of Gaussian distributions
Gaussian distributions are ubiquitous in modern science, and there are numerous ways to observe and understand their occurrence. They are, among others, the maximum entropy at fixed variance or the eigenvector of the Fourier transform. One very nice geometric characterization of multivariate Gaussian distributions is due to Maxwell [13, 4]. His theorem states that a random vector of dimension two or more has independent entries and is rotationally invariant if and only if its components are Gaussian, centered, with the same variance. In other words, such a random vector has a law with a density of the form
$H\mapsto\alpha\exp(-|H|_{2}^{2}/\gamma),$
@@
$H\mapsto\exp\left(-a\mathrm{Tr}(H^{2})+b\mathrm{Tr}(H)+c\right),$
@@
Theorem (Main).
For any $p\geq 1$ and $N\geq 2$, a real symmetric (resp. hermitian, resp. self-dual hermitian) random tensor $H$ of order $p$ and dimension $N$ (1.2) has :
    independent entries (up to symmetries),
    a law invariant by orthogonal (resp. unitary, resp. symplectic) transformation (Definition 2),
if and only if it belongs to the Gaussian Orthogonal (resp. Unitary, resp. Symplectic) Tensor Ensemble (1.3), that is it has a law with a density of the form
where $(\alpha,\beta,\gamma)\in\mathbb{R}^{*}_{+}\times\mathbb{R}\times\mathbb{R}^{*}_{+}$, and $\mathcal{I}$ is the symmetric tensor identity (Definition 1).
@@ 1.2 Real and complex tensors
A tensor is an element of the vector space $(\mathbb{C}^{N})^{\otimes p}$ for some $p\geq 1$ and $N\geq 1$. The integer $p$ is called the order, when $p=1$ (resp. $p=2$) we recover complex vectors (resp. matrices) space. The integer $N$ is called the dimension, note that we are dealing in this work only with cubic tensors, that is when each one of the $p$ legs has the same dimension $N$, because we focus on global invariance and not local one. For a complex tensor $H\in(\mathbb{C}^{N})^{\otimes p}$, the Frobenius norm of $H$ is the sum of the square of the modulus of the entries of the tensor, that is
$\|H\|_{\mathrm{F}}^{2}=\sum_{i_{1},\ldots,i_{p}}|H_{i_{1},\ldots,i_{p}}|^{2}.$
@@
Real symmetric and antisymmetric tensors.
Let $H\in(\mathbb{R}^{N})^{\otimes p}$ be a real tensor of order $p$ and dimension $N$ for some $p\geq 1,N\geq 1$. The symmetric group $S_{p}$ acts on $(\mathbb{R}^{N})^{\otimes p}$ by permutation of the indices
$H^{\sigma}_{i_{1},\ldots,i_{p}}=H_{i_{\sigma(1)},\ldots,i_{\sigma(p)}},$
for $\sigma\in S_{p}$. A tensor is said symmetric if
$H=H^{\sigma},$
for all $\sigma\in S_{p}$. Note that when $p=1$, a vector is always symmetric. We denote $\mathcal{S}^{(p)}(N)$ the set of real symmetric tensors of order $p$ and dimension $N$. Moreover, we define an antisymmetric tensor as a real tensor satisfying
$H^{\tau}_{i_{1},\ldots,i_{p}}=\epsilon(\tau)H_{i_{1},\ldots,i_{p}},$
for all $\tau\in S_{p}$ with no fixed point. Similarly, we denote $\mathcal{A}^{(p)}(N)$ the set of real antisymmetric tensors of order $p$ and dimension $N$. We say $(i_{1},\ldots,i_{p})\sim(j_{1},\ldots,j_{p})$ if $\forall k,i_{k}=j_{\sigma(k)}$ for some $\sigma\in\mathrm{S}_{p}$ and we denote
$\mathfrak{P}=\{(i_{1},\ldots,i_{p})\in\{1,\ldots,N\}^{p}|\ \ \exists j_{1},\ldots,j_{p/2}\text{ s.t. }(i_{1},\ldots,i_{p})\sim(j_{1},j_{1},\ldots,j_{p/2},j_{p/2})\},$
the set of paired indices. When $p$ is even, one can can check that if $H$ is antisymmetric, $H_{i_{1},\ldots,i_{p}}=0$ for all $(i_{1},\ldots,i_{p})\in\mathfrak{P}$. Denote $\Gamma_{i_{1},\ldots,i_{p}}$ the cardinality of the class of $(i_{1},\ldots,i_{p})$ under $\sim$, that is
$\Gamma_{i_{1},\ldots,i_{p}}=\frac{p!}{\prod_{j=1}^{N}c_{j}(i_{1},\ldots,i_{p})!},$
with $c_{j}(i_{1},\ldots,i_{p})$ the number of occurrences of $j$ in $(i_{1},\ldots,i_{p})$.
Definition 1.
@@
$\mathcal{I}_{i_{1},\ldots,i_{p}}=\mathds{1}_{(i_{1},\ldots,i_{p})\in\mathfrak{P}}/\Gamma_{i_{1},\ldots,i_{p}}\quad\quad\mbox{if }p\mbox{ even.}$.
Hermitian tensors.
Now, let $p$ be an even integer. Let $H\in(\mathbb{C}^{N})^{\otimes p}$ be a complex tensor of order $p$ and dimension $N$. We say that $H$ is hermitian if
$H=H^{(0)}+iH^{(1)},$
with $H^{(0)}\in\mathcal{S}^{(p)}(N)$ and $H^{(1)}\in\mathcal{A}^{(p)}(N)$. Hence, for any $(i_{1},\ldots,i_{p})\in\mathfrak{P}$, we have in particular $H_{i_{1},\ldots,i_{p}}=H^{(0)}_{i_{1},\ldots,i_{p}}\in\mathbb{R}$. We denote $\mathcal{H}^{(p)}(N)$ the set of Hermitian tensors of order $p$ and dimension $N$.
Self-dual hermitian tensors.
Let $p\equiv 2\pmod{4}$ be an integer and let $H\in(\mathbb{C}^{2N})^{\otimes p}$ be a complex tensor of order $p$ and dimension $2N$. The quaternion elements, denoted
$\mathbf{1}=\mathbf{e_{0}}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\quad\mathrm{and}\quad\mathbf{e_{1}}=\begin{pmatrix}i&0\\ 0&-i\end{pmatrix}\quad\mathrm{and}\quad\mathbf{e_{2}}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\quad\mathrm{and}\quad\mathbf{e_{3}}=\begin{pmatrix}0&-i\\ -i&0\end{pmatrix},$
@@
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\frac{a+d}{2}\mathbf{1}-i\frac{a-d}{2}\mathbf{e_{1}}-\frac{b-c}{2}\mathbf{e_{2}}+i\frac{b+c}{2}\mathbf{e_{3}}.$
Thus, identifying $(\mathbb{C}^{2N})^{\otimes p}$ with $(\mathbb{C}^{2}\otimes\mathbb{C}^{2})^{\otimes p/2}\times(\mathbb{C}^{N})^{\otimes p}$, the complex tensor $H$ may be canonically written as a tensor $Q$ of order $p$, dimension $N$ and with entries in $(\mathbb{C}^{2}\otimes\mathbb{C}^{2})^{\otimes p/2}$, given by
$Q=\sum_{\epsilon_{1},\ldots,\epsilon_{p/2}\in\{0,1,2,3\}}Q^{(\epsilon_{1},\ldots,\epsilon_{p/2})}\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}},$
with $Q^{(\epsilon_{1},\ldots,\epsilon_{p/2})}\in(\mathbb{C}^{N})^{\otimes p}$. Denote $\epsilon=(\epsilon_{1},\ldots,\epsilon_{p/2})$ and for $0\leq k\leq 3$, $n_{k}(\epsilon):=|\{s:\epsilon_{s}=k\}|$. Let $\mathcal{E}$ be the set of tuples $\epsilon$ such that $n_{1}(\epsilon)+n_{3}(\epsilon)\equiv n_{2}(\epsilon)\pmod{2}$, where the number of imaginary and the number of antisymmetric quaternions have same parity. We say that $Q$ (and then $H$) is self-dual hermitian if
$Q^{(\epsilon)}\in\mathcal{S}^{(p)}(N)\text{ if }\epsilon\in\mathcal{E}\quad\mathrm{and}\quad Q^{(\epsilon)}\in\mathcal{A}^{(p)}(N)\text{ otherwise}.$
@@
Similarity transformations.
Les $U$ be a nonsingular $N\times N$ matrix. For $p\geq 1$ and $N\geq 1$, the multilinear transformation of a tensor $H\in(\mathbb{C}^{N})^{\otimes p}$ under $U$ is a tensor denoted $U\cdot H\in(\mathbb{C}^{N})^{\otimes p}$. Three particular similarity transformations will capture our attention : the cases where $U$ belongs to $\mathrm{O}(N)$ the orthogonal group ($U^{T}U=I_{N}$), to $\mathrm{U}(N)$ the unitary group ($U^{*}U=I_{N}$) and to $\mathrm{Sp}(2N)$ the symplectic group ($U^{T}JU=J$ with $J=\mathbf{e_{2}}I^{\mathbb{H}}_{N}$ in symplectic notation). In these cases, the similarity transformations are given by
@@
    $(U\cdot H)_{i_{1},\ldots,i_{p}}=\sum_{j_{1},\ldots,j_{p}}H_{j_{1},\ldots,j_{p}}\prod_{t=1}^{p}U_{j_{t}i_{t}},$
@@
    $(U\cdot H)_{i_{1},\ldots,i_{p}}=\sum_{j_{1},\ldots,j_{p}}H_{j_{1},\ldots,j_{p}}\prod_{t=1}^{p/2}U_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}\bar{U}_{j_{2t}i_{2t}},$
@@
    $(U\cdot H)_{i_{1},\ldots,i_{p}}=\sum_{j_{1},\ldots,j_{p}}H_{j_{1},\ldots,j_{p}}\prod_{t=1}^{p/2}U_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}(-JUJ)_{j_{2t}i_{2t}}.$
Definition 2.
A real symmetric (resp. hermitian, resp. self-dual hermitian) random tensor $H$ is said orthogonal invariant (resp. unitary invariant, resp. symplectic invariant) if $H$ has the same law as $U\cdot H$ for any $U\in\mathrm{O}(N)$ (resp. $\mathrm{U}(N)$, resp. $\mathrm{Sp}(2N)$).
Remark 1.
In the complex case, one could also consider the transformation $(H\cdot(U^{(1)},\ldots,U^{(p/2)}))_{i_{1},\ldots,i_{p}}=\sum_{j_{1},\ldots,j_{p}}H_{j_{1},\ldots,j_{p}}\prod_{t=1}^{p/2}U^{(t)}_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}\bar{U}^{(t)}_{j_{2t}i_{2t}}$ where we act with a different unitary on each leg of the tensor. In this case we speak of local unitary invariance and many applications exist in quantum physics. We refer the reader to the rich works of Gurau [7], Lionni [12] or Sasakura [17] for instance.
Remark 2.
In the symplectic case, it is also possible to consider the similarity transformation $(U\cdot H)_{i_{1},\ldots,i_{p}}=\sum_{j_{1},\ldots,j_{p}}H_{j_{1},\ldots,j_{p}}\prod_{t=1}^{p}U_{j_{t}i_{t}}$, and in this case the trace invariants (see Section 1.4) are multigraphs where edges are decorated by a $J$ instead of multigraphs where odd halfedges are matched with even ones. We do this choice because we want to recover similarity transformations of the type $H\mapsto U^{-1}HU$ in the matrix case.

### v2

@@
[…]
Characterization of Gaussian Tensor Ensembles
[…]
Characterization of Gaussian Tensor Ensembles
[…]
Rémi BONNIN
[…]
R. Bonnin
[…]
Received July 09, 2025, in final form March 13, 2026; Published online March 31, 2026
[…]
The starting point of this work is a theorem due to Maxwell characterizing the distribution of a Gaussian vector with at least two coordinates. We define the Gaussian orthogonal, unitary and symplectic tensor ensembles for notions of real symmetric, hermitian and self-dual hermitian tensors which recover the classical vector and matrix Gaussian ensembles when the order is one and two. We give a complete family of invariant polynomials for orthogonal, unitary and symplectic transformations and we prove a Maxwell-type theorem for these Gaussian tensor distributions unifying and extending the ones known for vectors and matrices.
[…]
random tensors; Gaussian ensembles; Maxwell’s theorem; orthogonal/unitary/symplectic invariance
[…]
60B20; 81T32
@@ 1.1 Characterization of Gaussian distributions
Gaussian distributions are ubiquitous in modern science, and there are numerous ways to observe and understand their occurrence. They are, among others, the maximum entropy at fixed variance or the eigenvector of the Fourier transform. One very nice geometric characterization of multivariate Gaussian distributions is due to Maxwell [4, 13]. His theorem states that a¬random vector of dimension two or more has independent entries and is rotationally invariant if and only if its components are Gaussian, centered, with the same variance. In other words, such a random vector has a law with a density of the form
$H\mapsto\alpha\exp\bigl(-|H|_{2}^{2}/\gamma\bigr),$
@@
$H\mapsto\exp\bigl(-a\operatorname{Tr}\bigl(H^{2}\bigr)+b\operatorname{Tr}(H)+c\bigr),$
@@
{theorem*}
[main] For any $p\geq 1$ and $N\geq 2$, a real symmetric ((resp. hermitian, resp. self-dual hermitian)) random tensor $H$ of order $p$ and dimension $N$ (1.2) has
  • independent entries ((up to symmetries)),
  • a law invariant by orthogonal ((resp. unitary, resp. symplectic)) transformation ((see Definition 1.3)),
if and only if it belongs to the Gaussian orthogonal ((resp. unitary, resp. symplectic)) tensor ensemble (1.3), that is, it has a law with a density of the form
$H\mapsto\alpha\exp\bigl(-\|H-\beta\mathcal{I}\|_{\mathrm{F}}^{2}/\gamma\bigr),$
where $(\alpha,\beta,\gamma)\in\mathbb{R}^{*}_{+}\times\mathbb{R}\times\mathbb{R}^{*}_{+}$, and $\mathcal{I}$ is the symmetric tensor identity ((Definition 1.1)).
@@ 1.2 Real and complex tensors
A tensor is an element of the vector space $\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$ for some $p\geq 1$ and $N\geq 1$. The integer $p$ is called the order, when $p=1$ (resp. $p=2$) we recover complex vectors (resp. matrices) space. The integer $N$ is called the dimension, note that we are dealing in this work only with cubic tensors, that is, when each one of the $p$ legs has the same dimension $N$, because we focus on global invariance and not local one. For a complex tensor $H\in\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$, the Frobenius norm of $H$ is the sum of the square of the modulus of the entries of the tensor, that is,
$\|H\|_{\mathrm{F}}^{2}=\sum_{i_{1},\dots,i_{p}}|H_{i_{1},\dots,i_{p}}|^{2}.$
@@
Real symmetric and antisymmetric tensors. Let $H\in\bigl(\mathbb{R}^{N}\bigr)^{\otimes p}$ be a real tensor of order $p$ and dimension $N$ for some $p\geq 1$, $N\geq 1$. The symmetric group $S_{p}$ acts on $\bigl(\mathbb{R}^{N}\bigr)^{\otimes p}$ by permutation of the indices $H^{\sigma}_{i_{1},\dots,i_{p}}=H_{i_{\sigma(1)},\dots,i_{\sigma(p)}}$, for $\sigma\in S_{p}$. A tensor is said symmetric if $H=H^{\sigma}$, for all $\sigma\in S_{p}$. Note that when $p=1$, a vector is always symmetric. We denote $\mathcal{S}^{(p)}(N)$ the set of real symmetric tensors of order $p$ and dimension $N$. Moreover, we define an antisymmetric tensor as a real tensor satisfying $H^{\tau}_{i_{1},\dots,i_{p}}=\epsilon(\tau)H_{i_{1},\dots,i_{p}}$, for all $\tau\in S_{p}$. Similarly, we denote $\mathcal{A}^{(p)}(N)$ the set of real antisymmetric tensors of order $p$ and dimension $N$. We say $(i_{1},\dots,i_{p})\sim(j_{1},\dots,j_{p})$ if $\forall k,i_{k}=j_{\sigma(k)}$ for some $\sigma\in\mathrm{S}_{p}$ and we denote
$\mathfrak{P}=\{(i_{1},\dots,i_{p})\in\{1,\dots,N\}^{p}\mid\exists j_{1},\dots,j_{p/2}\text{ s.t.\ }(i_{1},\dots,i_{p})\sim(j_{1},j_{1},\dots,j_{p/2},j_{p/2})\},$
the set of paired indices. When $p$ is even, one can check that if $H$ is antisymmetric, $H_{i_{1},\dots,i_{p}}=0$ for all $(i_{1},\dots,i_{p})\in\mathfrak{P}$. Denote $\Gamma_{i_{1},\dots,i_{p}}$ the cardinality of the class of $(i_{1},\dots,i_{p})$ under $\sim$, that is,
$\Gamma_{i_{1},\dots,i_{p}}=\frac{p!}{\prod_{j=1}^{N}c_{j}(i_{1},\dots,i_{p})!},$
with $c_{j}(i_{1},\dots,i_{p})$ the number of occurrences of $j$ in $(i_{1},\dots,i_{p})$.
Definition 1.1.
@@
$\mathcal{I}_{i_{1},\dots,i_{p}}=\mathds{1}_{(i_{1},\dots,i_{p})\in\mathfrak{P}}/\Gamma_{i_{1},\dots,i_{p}}\qquad\mbox{if }p\mbox{ even.}$
Hermitian tensors. Now, let $p$ be an even integer. Let $H\in\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$ be a complex tensor of order $p$ and dimension $N$. We say that $H$ is hermitian if $H=H^{(0)}+iH^{(1)}$, with $H^{(0)}\hskip-0.5pt\in\hskip-0.5pt\mathcal{S}^{(p)}(N)$ and $H^{(1)}\hskip-1.0pt\in\hskip-0.5pt\mathcal{A}^{(p)}(N)$. Hence, for any $(i_{1},\dots,i_{p})\hskip-0.5pt\in\hskip-0.5pt\mathfrak{P}$, we have in particular $H_{i_{1},\dots,i_{p}}=H^{(0)}_{i_{1},\dots,i_{p}}\hskip-1.0pt\in\hskip-0.5pt\mathbb{R}$. We denote $\mathcal{H}^{(p)}(N)$ the set of Hermitian tensors of order $p$ and dimension $N$.
Self-dual hermitian tensors. Let $p\equiv 2\pmod{4}$ be an integer and let $H\in\bigl(\mathbb{C}^{2N}\bigr)^{\otimes p}$ be a complex tensor of order $p$ and dimension $2N$. The quaternion elements, denoted
$\mathbf{1}=\mathbf{e_{0}}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad\mathbf{e_{1}}=\begin{pmatrix}{\rm i}&0\\ 0&-{\rm i}\end{pmatrix},\qquad\mathbf{e_{2}}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\qquad\mathbf{e_{3}}=\begin{pmatrix}0&-{\rm i}\\ -{\rm i}&0\end{pmatrix}$,
@@
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\frac{a+d}{2}\mathbf{1}-{\rm i}\frac{a-d}{2}\mathbf{e_{1}}-\frac{b-c}{2}\mathbf{e_{2}}+{\rm i}\frac{b+c}{2}\mathbf{e_{3}}.$
Thus, identifying $\bigl(\mathbb{C}^{2N}\bigr)^{\otimes p}$ with $\bigl(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\bigr)^{\otimes p/2}\times\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$, the complex tensor $H$ may be canonically written as a tensor $Q$ of order $p$, dimension $N$ and with entries in $\bigl(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\bigr)^{\otimes p/2}$, given by
$Q=\sum_{\epsilon_{1},\dots,\epsilon_{p/2}\in\{0,1,2,3\}}Q^{(\epsilon_{1},\dots,\epsilon_{p/2})}\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}},$
with $Q^{(\epsilon_{1},\dots,\epsilon_{p/2})}\in\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$. Denote $\epsilon=(\epsilon_{1},\dots,\epsilon_{p/2})$ and for $0\leq k\leq 3$, $n_{k}(\epsilon):=|\{s\colon\epsilon_{s}=k\}|$. Let $\mathcal{E}$ be the set of tuples $\epsilon$ such that $n_{1}(\epsilon)+n_{3}(\epsilon)\equiv n_{2}(\epsilon)\pmod{2}$, where the number of imaginary and the number of antisymmetric quaternions have same parity. We say that $Q$ (and then $H$) is self-dual hermitian if
[…]
@@
We briefly discuss the motivation from the perspective of theoretical physics. Rephrasing Dyson and Mehta, self-dual Hermitian operators correspond to Hamiltonians of systems (with odd spin) that are invariant under time reversal but not rotationally invariant. More precisely, a quaternion element
$q=q^{(0)}\mathbf{1}+q^{(1)}\mathbf{e_{1}}+q^{(2)}\mathbf{e_{2}}+q^{(3)}\mathbf{e_{3}}$
is called real if $q^{(0)},q^{(1)},q^{(2)},q^{(3)}\in\mathbb{R}$ (note that this differs from $q\in\mathbb{R}^{2\times 2}$). If the quaternion is complex, its quaternion conjugate is
$\bar{q}=q^{(0)}\mathbf{1}-q^{(1)}\mathbf{e_{1}}-q^{(2)}\mathbf{e_{2}}-q^{(3)}\mathbf{e_{3}},$
and its Hermitian conjugate is
$q^{\dagger}=q^{(0)*}\mathbf{1}-q^{(1)*}\mathbf{e_{1}}-q^{(2)*}\mathbf{e_{2}}-q^{(3)*}\mathbf{e_{3}}.$
For $H$ a matrix of dimension $2N$ written as $Q=(q_{ij})$ of dimension $N$ with quaternionic entries, time reversal invariance and Hermiticity require respectively $\bar{q}_{ij}=q_{ji}$ and $q_{ij}^{\dagger}=q_{ji}$, which gives $Q^{(0)}:=\bigl(q_{ij}^{(0)}\bigr)_{i,j}$ is real symmetric and $Q^{(1)}$, $Q^{(2)}$, $Q^{(3)}$ are real antisymmetric. We observe here that the time invariance forces $Q^{(2)}$ to be antisymmetric because $\mathbf{e_{2}}$ is the only antisymmetric quaternion basis element and the Hermiticity forces $Q^{(1)}$, $Q^{(3)}$ to be antisymmetric because $\mathbf{e_{1}}$ and $\mathbf{e_{3}}$ are the complex quaternion basis elements. This is the first motivation of our definition of self-dual hermitian tensors. For a tensor ($p$-spin) model, we require $Q^{(\epsilon)}$ to be real symmetric or antisymmetric according to whether the number of antisymmetric quaternions factors $\mathbf{e_{2}}$ appearing in the decomposition $\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}$ has the same parity as the number of complex quaternions factors $\mathbf{e_{1}}$ and $\mathbf{e_{3}}$, or not.
Example 1.2.
For $p=2$, this coincides with the classical definition of self-dual Hermitian matrix. In this case,
$Q=Q^{(0)}\otimes\mathbf{1}+Q^{(1)}\otimes\mathbf{e_{1}}+Q^{(2)}\otimes\mathbf{e_{2}}+Q^{(3)}\otimes\mathbf{e_{3}},$
with $Q^{(0)}=\bigl(Q^{(0)}\bigr)^{\mathsf{T}}$ and $Q^{(k)}=-\bigl(Q^{(k)}\bigr)^{\mathsf{T}}$ for $k\in\{1,2,3\}$.
The first genuine tensor example is $p=6$. Here we have $p/2=3$ and for $\epsilon=(\epsilon_{1},\epsilon_{2},\epsilon_{3})$, the parity condition is $n_{1}(\epsilon)+n_{3}(\epsilon)\equiv n_{2}(\epsilon)\pmod{2}$, or in other words $\epsilon\in\mathcal{E}$ if the number of non-zero coefficients is odd. We have, for example,
$(0,0,0),(1,1,0),(1,3,0)\in\mathcal{E},\qquad\text{but }(1,0,0),(1,1,1),(1,3,2)\notin\mathcal{E},$
meaning that $Q^{(0,0,0)}$, $Q^{(1,1,0)}$, $Q^{(1,3,0)}$ are symmetric tensors of order $3$ and $Q^{(1,0,0)}$, $Q^{(1,1,1)}$, $Q^{(1,3,2)}$ are antisymmetric ones.
Similarity transformations. Let $U$ be a nonsingular $N\times N$ matrix. For $p\geq 1$ and $N\geq 1$}, the multilinear transformation of a tensor $H\in\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$ under $U$ is a tensor denoted $U\cdot H\in\bigl(\mathbb{C}^{N}\bigr)^{\otimes p}$. Three particular similarity transformations will capture our attention: the cases where $U$ belongs to $\mathrm{O}(N)$ the orthogonal group $\bigl(U^{\mathsf{T}}U=I_{N}\bigr)$, to $\mathrm{U}(N)$ the unitary group ($U^{*}U=I_{N}$}) and to $\mathrm{Sp}(2N)$ the symplectic group ($U^{\mathsf{T}}JU=J$} with $J=\mathbf{e_{2}}I^{\mathbb{H}}_{N}$ in symplectic notation). In these cases, the similarity transformations are given by
@@
    $(U\cdot H)_{i_{1},\dots,i_{p}}=\sum_{j_{1},\dots,j_{p}}H_{j_{1},\dots,j_{p}}\prod_{t=1}^{p}U_{j_{t}i_{t}},$
@@
    $(U\cdot H)_{i_{1},\dots,i_{p}}=\sum_{j_{1},\dots,j_{p}}H_{j_{1},\dots,j_{p}}\prod_{t=1}^{p/2}U_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}\bar{U}_{j_{2t}i_{2t}},$
@@
    $(U\cdot H)_{i_{1},\dots,i_{p}}=\sum_{j_{1},\dots,j_{p}}H_{j_{1},\dots,j_{p}}\prod_{t=1}^{p/2}U_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}(-JUJ)_{j_{2t}i_{2t}}.$
Definition 1.3.
A real symmetric ((resp. hermitian, resp. self-dual hermitian)) random tensor $H$ is said orthogonal invariant ((resp. unitary invariant, resp. symplectic invariant)) if $H$ has the same law as $U\cdot H$ for any $U\in\mathrm{O}(N)$ ((resp. $\mathrm{U}(N)$, resp. $\mathrm{Sp}(2N))$.
Remark 1.4.
In the complex case, one could also consider the transformation
$\bigl(H\cdot\bigl(U^{(1)},\dots,U^{(p/2)}\bigr)\bigr)_{i_{1},\dots,i_{p}}=\sum_{j_{1},\dots,j_{p}}H_{j_{1},\dots,j_{p}}\prod_{t=1}^{p/2}U^{(t)}_{j_{2t-1}i_{2t-1}}\prod_{t=1}^{p/2}\bar{U}^{(t)}_{j_{2t}i_{2t}}$,
where we act with a different unitary on each leg of the tensor. In this case we speak of local unitary invariance and many applications exist in quantum physics. We refer the reader to the rich works of Gurau [7], Lionni [12] or Sasakura [17] for instance.
Remark 1.5.
In the symplectic case, it is also possible to consider the similarity transformation $(U\cdot H)_{i_{1},\dots,i_{p}}=\sum_{j_{1},\dots,j_{p}}H_{j_{1},\dots,j_{p}}\prod_{t=1}^{p}U_{j_{t}i_{t}}$}, and in this case the trace invariants (see Section 1.4) are multigraphs where edges are decorated by a $J$ instead of multigraphs where odd halfedges are matched with even ones. We do this choice because we want to recover similarity transformations of the type $H\mapsto U^{-1}HU$ in the matrix case.